OpenStudy (anonymous):

If log3 (x - 5) = 2 What is ln x? Please show work

OpenStudy (anonymous):

\[\log _{3} (x - 5) = 2\]

OpenStudy (anonymous):

x=14

OpenStudy (lgbasallote):

\[\log_3 (x-5) = 2\] \[3^2 = x - 5\] \[9 = x - 5\] \[9+5 = x\] \[14 = x\] \[\ln x = \ln (14)\]

OpenStudy (anonymous):

14